The main constituents of air are 80% of nitrogen molecules of molar mass 28 kg/k mol and
20% oxygen molecules of molar mass 32 kg/k mol. The mass of air in 50 cm3 volume at
9.3 × 104 Pa pressure at room temperature (20°C) is
[use ideal gas law] (R = 8.314 × 103 J/k mol/K)
(1) 5.5 kg (2) 55 × 10–3 kg
(3) 55 kg (4) 5.5 × 10–5 kg
1
Expert's answer
2019-04-24T10:23:36-0400

Dalton's law of partial pressure

"p=p_1+p_2"

From ideal gas law we have 3 equation

"p_1V=\frac{m_1}{M_1}RT""p_2V=\frac{m_2}{M_2}RT""pV=(\frac{m_1}{M_1}+\frac{m_2}{M_2})RT"

But,

"m_1=0.2m; m_2=0.8m""pV=(\frac{0.2m}{M_1}+\frac{0.8m}{M_2})RT""m=\frac{pV}{(\frac{0.8}{M_1}+\frac{0.2}{M_2})RT}""m=\frac{93000\cdot 0.00005}{(\frac{0.8}{28}+\frac{0.2}{32})8314\cdot 293}=5.5\cdot 10^{-5} kg"

Answer: (4) 5.5*10^-5